By selling a pen for rs.165, mukti gains 10%. If she had sold it for rs.157.50, what percentage would she have lost or gained?
Answer by mananth(16949)

if 110 is the SP CP=100
165 was the SP
110 ------100
165-------x
x/100=165/110
x=(100*165)/110
CP= 150
sold for 157.50

150--------------7.50
100--------------x
x/7.50 = 100/150
x=7.50*100/150
She makes a profit of 5%